OpenStudy (anonymous):

e^x / 1 - e^x = 2 find the x

OpenStudy (mertsj):

\[\frac{e^x}{1-e^x}=2\] \[2-2e^x=e^x \] \[2=3e^x\] \[\ln 2=\ln 3e^x\] \[\ln 2=\ln 3+\ln e^x\] \[\ln 2-\ln 3=x \ln e=x\]